# The Assouad dimension of Kakeya sets in $\mathbb{R}^{3}$

### v1

@@ 6 Kakeya estimates for tubes satisfying the Convex Wolff Axioms at all scales
Stickiness is a multi-scale property, in the following sense: if $\mathbb{T}$ is sticky, then for each $\rho\geq\delta$ we can cover $\mathbb{T}$ by a sticky collection of $\rho$-tubes. This multi-scale property was exploited throughout the arguments in [16]. Theorem 5.2 differs from [16, Theorem 1.1], because the stickiness hypothesis has been replaced by the hypothesis that the tubes satisfy the Convex Wolff axioms at every scale. As we will see below, however, this latter hypothesis is sufficient in order to repeat the arguments from [16]. Indeed, with two exceptions that will be explained below, the arguments from [16] can be repeated mutatis mutatis to obtain Theorem 5.2.
The proof of [16, Theorem 1.1] is divided into several sections. The paper begins by supposing that the result is false, and studying the structure of a (hypothetical) counter-example. In Section 2, the authors define what it means for a pair $(\mathbb{T},Y)$ of tubes and their associated shading to be a worst possible counter-example to [16, Theorem 1.1]; this is called an $\varepsilon$-extremal pair. In our discussion below, we will call it a $\varepsilon$-extremal pair. In [16], Section 3, it is shown that an $\varepsilon$-extremal pair $(\mathbb{T},Y)$ must look coarsely self similar at every intermediate scale $\rho\in[\delta,1]$. More precisely, for every such $\rho$ it is possible to cover the tubes in $\mathbb{T}$ by a collection of $\rho$-tubes, so that the corresponding collection of $\rho$-tubes is a worst possible counter-example to [16, Theorem 1.1] at scale $\rho$, and the $\delta$-tubes inside each $\rho$-tube also a form a (re-scaled) worst-possible counter-example at scale $\delta/\rho$. The analogous statement and proof in our setting are nearly identical; we will state the former and briefly sketch the latter.
@@
Finally, in Section 7 of [16], the authors recall an analogue of Wolff’s circular maximal function estimate [19, 14], and they show how this estimate contradicts the assumption that there exits a collection of sticky tubes whose union has volume much smaller than 1, thereby completing the proof of the theorem. This section of [16] uses the assumption that the tubes point in different directions, in order to obtain a rich planar arrangement of curves to which Wolff’s circular maximal theorem can be applied. With some modifications, the direction-separated assumption can be replaced with a weaker ball condition; we will explain this in detail in Sections 6.5 and 7.
@@ 6.1 Relevant definitions from [16]
Before describing the arguments in [16], it will be helpful to recall a few key definitions. First, a minor technical annoyance: tubes and cubes in [16] are defined slightly differently than in the present paper. In [16], a $\delta$-tube in $\mathbb{R}^{3}$ is defined to be a set of the form $N_{6\delta}(\ell)\cap[-1,1]^{3}$, where $\ell$ is a line in $\mathbb{R}^{3}$. This definition is adopted for technical reasons, and for consistency with the results in [16] we will use this definition throughout Section (6); the distinction between this definition and our previous definition in Section 1 is not important in the arguments that follow. Second, in [16], a shading $Y(T)\subset T$ is defined to be a union of axis-aligned $\delta$-cubes contained in $T$. Again, the distinction between this definition and our previous one from Section 2.1 is harmless: replace each tube $T$ (in the sense of Section 1) and a each shading $Y(T)$ (in the sense of Section 2.1) with the corresponding tube in the sense of [16] and the the shadings consisting of the union of axis-aligned $\delta$-cubes $Q$ that satisfy $|Q\cap Y(T)|\geq\frac{1}{100}|Y(T)|/|T|$. This gives us a collection of tubes and shadings in the sense of [16], and if the conclusion of Theorem 1.5 holds for this latter collection of tubes and shadings, then it also holds for the original collection of tubes and shadings, except Inequality (1.1) has been weakened by an additional (harmless) factor of $\delta^{\eta}$.
@@
We will often refer to a collection of $\delta$-tubes $\mathbb{T}$, and a shading $Y(T)$ of these tubes. We will refer to this pair as $(\mathbb{T},Y)_{\delta}$. We use $E_{\mathbb{T}}$ to denote the set $\bigcup_{T\in\mathbb{T}}Y(T)$; the shading $Y$ will always be apparent from context. For $p\in\mathbb{R}^{3}$, we define
@@
$|\operatorname{dir}(T)\cdot V(p)|\leq\delta\quad\textrm{for all}\ (T,p)\in\mathbb{T}\times\mathbb{R}^{3}\ \textrm{with}\ p\in Y(T),$
@@ 6.2 Extremal collection of tubes, and multi-scale structure
$\Big{|}\bigcup_{T\in\mathbb{T}}Y(T)\Big{|}\leq\delta^{\sigma^{\prime}}.$ (6.2)
@@
Let $\varepsilon,\delta>0$. We say a pair $(\mathbb{T},Y)_{\delta}$ is $\varepsilon$-extremal if
@@
    $\Big{|}\bigcup_{T\in\mathbb{T}}Y(T)\Big{|}\leq\delta^{\sigma-\varepsilon}.$
It is immediate from the above definitions that for every $\varepsilon,\delta_{0}>0$, there exists a $\varepsilon$-extremal collection of tubes $(\mathbb{T},Y)_{\delta}$ for some $\delta\in(0,\delta_{0}]$; cf. [16, Lemma 3.1].
@@
For all $\varepsilon>0$, there exists $\eta>0$ and $\delta_{0}>0$ so that the following holds for all $\delta\in(0,\delta_{0}]$. Let $(\mathbb{T},Y)_{\delta}$ be a $\eta$-extremal collection of tubes, and let $\rho\in[\delta^{1-\varepsilon},\delta^{\varepsilon}]$. Then there is a refinement $(\mathbb{T}^{\prime},Y^{\prime})_{\delta}$ of $(\mathbb{T},Y)_{\delta}$ and a balanced cover $(\tilde{\mathbb{T}},\tilde{Y})_{\rho}$ of $(\mathbb{T}^{\prime},Y^{\prime})_{\delta}$ with the following properties.
@@
The proof is nearly identical to the proof of the corresponding statement in [16, Proposition 3.2]. In [16], the first major step was to establish the existence of a cover $\mathbb{T}_{\rho}$ of $\mathbb{T}$ (or more accurately, a large subset $\mathbb{T}^{\prime}\subset\mathbb{T}$), with the property that both $\mathbb{T}_{\rho}$ and the the re-scaled sets $\mathbb{T}^{\prime}[T_{\rho}]$ are sticky. In our case, this step was handled by Lemma 2.13. The second major step was to establish a shading on $\mathbb{T}_{\rho}$ and a refinement of the shadings on the sets $\mathbb{T}^{\prime}[T_{\rho}]$ so that not too many tubes pass through each point (and as a consequence, the corresponding collections are $\varepsilon$-extremal).
@@ 6.3 Local and global grains
For all $\varepsilon,\delta_{0}>0$, there exists $\delta\in(0,\delta_{0}]$ and a $\varepsilon$-extremal set of tubes $(\mathbb{T},Y)_{\delta}$ with the following two properties.
@@
    There is a 1-Liptschitz function $f:[-1,1]\rightarrow\mathbb{R}$ so that $(E_{\mathbb{T}}\cap\{z=z_{0}\})\cdot(1,f(z_{0}),0)$ is a $(\delta,1-\sigma,\delta^{-\varepsilon})_{1}$-ADset for each $z_{0}\in[-1,1]$.
@@
For all $\varepsilon,\delta_{0}>0$, there exists $\delta\in(0,\delta_{0}]$ and a $\varepsilon$-extremal set of tubes $(\mathbb{T},Y)_{\delta}$ with the following properties.
  1. 1’
@@ 6.4 The slope function has large slope
For all $\varepsilon,\delta_{0}>0$, there exists $\delta\in(0,\delta_{0}]$ and a $\varepsilon$-extremal set of tubes $(\mathbb{T},Y)_{\delta}$ with the following properties:
  1. 1”
@@
and that $(\mathbb{R}^{2}\times J)\cap\bigcup_{T\in\mathbb{T}}Y(T)$ can be covered by a union of $\leq\delta^{-\eta}|J|^{-2+\sigma}$ cubes of side-length $J$.
@@
Here is a summary of Steps 1-3. After a refinement, we may suppose that for each point $p\in F:=(\mathbb{R}^{2}\times J)\cap\bigcup_{T\in\mathbb{T}}Y(T)$,
@@
After a translation and a rotation, we may suppose that $J=[0,\rho^{1/2}]$ and $f(0)=0$. Since $|f^{\prime\prime}(z)|\lesssim 1$, if $|f^{\prime}(z)|\leq|J|$ for some $z\in J$, then $|f^{\prime}(z)|\lesssim|J|$ for all $z\in J$, and so $|f(z)|\lesssim|J|^{2}$. We identify a subset $F_{2}\subset F$ with $|F_{2}|\geq\delta^{4\eta}|F|$ so that
@@
    has a large intersection with $F_{2}$:
@@
This finishes the summary of Steps 1-3 and we have introduced the necessary geometric objects and notations.
@@
But or each $T\in\mathbb{T}_{S}$, we have
@@
By [16, Lemma 6.2], we have
@@
$\Sigma_{S}=\{(x,y,z)\in[0,1]^{3}:|(x-x_{S},0,z)\cdot\tilde{V}(p_{S})|\leq\delta^{-2C_{0}\eta}\rho^{1/2}.\}$
@@
#$\#\mathbb{T}_{S}\leq|\Sigma_{S}|\delta^{-\eta}(\#\mathbb{T})\leq\delta^{-(2C+1)\eta}\rho^{1/2}(\#\mathbb{T}).$
@@
The proof of Proposition 6.5 is the same as [16, Proposition 6.1] using Lemma 6.6 and [16, Lemma 6.4].
@@ 6.5 Twisted projections, and the completion of the proof
Recalling Definition 6.1 from [16], for $f\colon[-1,1]\to\mathbb{R}$ we define the twisted projection $\pi_{f}(x,y,z)=(x+f(z)y,z)$. The consequence of Proposition 6.5 is the following analogue of [16, Corollary 6.5]
@@
For all $\varepsilon,\delta_{0}>0$, there exists $\delta\in(0,\delta_{0}]$; a set pair $(\mathbb{T},Y)$ of $\delta$-tubes satisfying the Convex Wolff Axioms with error $\delta^{-\varepsilon}$ and an $\delta^{\varepsilon}$-dense shading; and a function $f\colon[-1,1]\to\mathbb{R}$ with $1\leq|f^{\prime}(z)|\leq 2$ and $|f^{\prime\prime}(z)|\leq 1/100$ for all $z\in[-1,1]$, so that
@@
To complete the proof of Theorem 5.2, it suffices to establish the following estimate, which is the analogue of [16, Proposition 7.1]
@@
Indeed, comparing Lemma 6.7 and Proposition 6.8, we conclude that $\sigma=0$, which completes the proof of Theorem 5.2. It remains to prove Proposition 6.8; we defer this to the next section.
@@ 7 Twisted projections of tubes satisfying the Convex Wolff Axioms
Our goal in this section is to prove Proposition 6.8. In fact, we will prove a slightly stronger statement. We say a set of tubes $\mathbb{T}$ satisfies the Tube Wolff Axioms with error $C$, if for every $\rho$-tube $T_{\rho}$, at most $C|T_{\rho}|(\#\mathbb{T})$ tubes from $\mathbb{T}$ are contained in $T_{\rho}$; since $\rho$-tubes are convex, this is a special case of the Convex Wolff Axioms. We will prove the following mild generalization of Proposition 6.8
@@
If $\mathbb{T}$ is a set of $\delta$-tubes contained in a bounded set such as $B(0,1)$, then we can identify the coaxial line of each tube with a point in $\mathbb{R}^{4}$. If $\mathbb{T}$ satisfies the Tube Wolff Axioms, then the corresponding set of points in $\mathbb{R}^{4}$ forms a $(\delta,2,O(1))$-Frostman set (technically the corresponding set of points might be a multi-set, but this is harmless—we can either extend the definition of a Orponen-Shmerkin-set to include this possibility, or infinitesimally perturb the tubes to avoid this situation).
@@
where each $B_{i}$ in the covering is a cube and $r(B_{i})$ is the side-length of $B_{i}$. By [3, Lemma 3.13], if $\mathcal{H}^{s}_{\infty}(A)=\kappa>0$, then there exists a $(\delta,s,100)$-Katz-Tao set $A^{\prime}\subset A$ with $\mathcal{E}_{\delta}(A^{\prime})\gtrsim\kappa\delta^{-s}$. Thus in order to prove the lemma, it suffices to obtain the bound
@@
Since $A$ is a $(\delta,s,C)$-Frostman set, $\mathcal{E}_{\delta}(A\cap B_{i})\leq C\lambda^{s}\mathcal{E}_{\delta}(A)$. The number of $B_{i}$ in the cover with $r(B_{i})\sim\lambda$ is $\gtrsim_{n,s}\lambda^{-s}|\log\delta|^{-1}C^{-1}$. This establishes (7.1) and concludes the proof. ∎
@@
Definition 7.7 ($\eta$-uniform set up to scale $\delta$).
@@
Let $\eta>0$ be a small parameter and $\delta\in 2^{-\eta^{-1}\mathbb{N}}$. Given an arbitrary set $A\subset[0,1]^{n}$, [13, Lemma 2.15] says there exists an $\eta$ uniform set $A^{\prime}\subset A$ up to scale $\delta$ and $\mathcal{E}_{\delta}(A^{\prime})\gtrsim|\log\delta|^{-\eta^{-1}}\mathcal{E}_{\delta}(A)$. We call $A^{\prime}$ an $\eta$-uniform refinement (usually $\delta$ is clear from context and we omit it from notation).
@@
Given an $\eta$-uniform set $A$ with $\delta$-covering number $\delta^{-s}$, the next lemma finds an Orponen-Shmerkin-set if one zooms in at sufficiently small scale.
@@
Start with the output of [15, Lemma 5.21] applying to $A$, if there exists $t_{j}\geq t_{j+1}$, apply [15, Lemma 5.20] to merge the two intervals $[a_{j},a_{j+1}]$ and $[a_{j+1},a_{j+2}]$ to $[a_{j},a_{j+2}]$ and the two numbers $t_{j},t_{j+1}$ becomes $t_{j}^{\prime}$. Rename the set of intervals and $t_{j}$s and replace $J$ by $J-1$. Iterate the process until $t_{j}<t_{j+1}$ for all $j$. Now we have obtained a decomposition of scales $\{[a_{j},a_{j+1}]\}_{j=1}^{J}$ such that
@@
Divide the unit cube $[0,1]^{n}$ into a union of cubes $\tilde{Q}$ of side length $\rho^{s/d}\leq\rho$. For each $\tilde{Q}$, let $A_{\tilde{Q}}$ denote the translation that maps the center of $Q_{0}$ to the center of $\tilde{Q}$. Define $A=\sqcup A_{\tilde{Q}}(A_{0})$. Since the side-length of $\tilde{Q}$ is greater than the side-length of $Q_{0}$, $A_{\tilde{Q}}(A_{0})$ are disjoint, and so $\mathcal{E}_{\delta}(A)\gtrsim\delta^{-s}$.
@@
$\mathcal{E}_{\delta}(A\cap B_{r})\lesssim C(\frac{r}{\rho^{s/n}})^{n}\cdot(\frac{\rho}{\delta})^{s}\lesssim C(\frac{r}{\delta})^{s}.\qed$
We prove a single scale version of Proposition 7.1.
@@
The quantity on the RHS of (7.2) will be closely related to the graph of cinematic functions (see [16, Definition 7.1]). To estimate it, we recall [16, Theorem 7.2].

### v2

@@ 6 Kakeya estimates for tubes satisfying the Convex Wolff Axioms at every scale
Stickiness is a multi-scale property, in the following sense: if $\mathbb{T}$ is sticky, then for each $\rho\geq\delta$ we can cover $\mathbb{T}$ by a sticky collection of $\rho$-tubes. This multi-scale property was exploited throughout the arguments in [16]. Theorem 5.2 differs from [16, Theorem 1.1], because the stickiness hypothesis has been replaced by the hypothesis that the tubes satisfy the Convex Wolff axioms at every scale. As we will see below, however, this latter hypothesis is sufficient in order to repeat the arguments from [16]. Indeed, with two exceptions that will be explained below, the arguments from [16] can be repeated mutatis mutandis to obtain Theorem 5.2.
The proof of [16, Theorem 1.1] is divided into several sections. The paper begins by supposing that the result is false, and studying the structure of a (hypothetical) counter-example. In Section 2, the authors define what it means for a pair $(\mathbb{T},Y)$ of tubes and their associated shading to be a worst possible counter-example to [16, Theorem 1.1]; this is called an $\varepsilon$-extremal pair, and we will adopt this terminology in our discussion below. In [16], Section 3, it is shown that an $\varepsilon$-extremal pair $(\mathbb{T},Y)$ must look coarsely self similar at every intermediate scale $\rho\in[\delta,1]$. More precisely, for every such $\rho$ it is possible to cover the tubes in $\mathbb{T}$ by a collection of $\rho$-tubes, so that the corresponding collection of $\rho$-tubes is a worst possible counter-example to [16, Theorem 1.1] at scale $\rho$, and the $\delta$-tubes inside each $\rho$-tube also form a (rescaled) worst-possible counter-example at scale $\delta/\rho$. The analogous statement and proof in our setting are nearly identical; we will state the former and briefly sketch the latter.
@@
Finally, in Section 7 of [16], the authors recall an analogue of Wolff’s circular maximal function estimate [19, 14], and they show how this estimate contradicts the assumption that there exists a collection of sticky tubes whose union has volume much smaller than 1, thereby completing the proof of the theorem. This section of [16] uses the assumption that the tubes point in different directions in order to obtain a rich planar arrangement of curves to which Wolff’s circular maximal theorem can be applied. With some modifications, the direction-separated assumption can be replaced with a weaker ball condition. We will explain this in detail in Sections 6.5 and 7.
@@ 6.1 Relevant definitions from [16]
Before describing the arguments in [16], it will be helpful to recall a few key definitions. First, a minor technical annoyance: tubes and cubes in [16] are defined slightly differently than in the present paper. In [16], a $\delta$-tube in $\mathbb{R}^{3}$ is defined to be a set of the form $N_{6\delta}(\ell)\cap[-1,1]^{3}$, where $\ell$ is a line in $\mathbb{R}^{3}$. This definition is adopted for technical reasons, and for consistency with the results in [16] we will use this definition throughout Section 6; the distinction between this definition and our previous definition in Section 1 is not important in the arguments that follow. Second, in [16], a shading $Y(T)\subset T$ is defined to be a union of axis-aligned $\delta$-cubes contained in $T$. Again, the distinction between this definition and our previous one from Section 2.1 is harmless: replace each tube $T$ (in the sense of Section 1) and a each shading $Y(T)$ (in the sense of Section 2.1) with the corresponding tube in the sense of [16] and the shadings consisting of the union of axis-aligned $\delta$-cubes $Q$ that satisfy $|Q\cap Y(T)|\geq\frac{1}{100}|Y(T)|/|T|$. This gives us a collection of tubes and shadings in the sense of [16], and if the conclusion of Theorem 1.5 holds for this latter collection of tubes and shadings, then it also holds for the original collection of tubes and shadings, except Inequality (1.1) has been weakened by an additional (harmless) factor of $1/1000$.
@@
We will often refer to a collection of $\delta$-tubes $\mathbb{T}$, and a shading $Y(T)$ of these tubes. We will refer to this pair as $(\mathbb{T},Y)_{\delta}$. We use $E_{\mathbb{T}}$ to denote the set $\bigcup_{\mathbb{T}}Y(T)$; the shading $Y$ will always be apparent from context. For $p\in\mathbb{R}^{3}$, we define
@@
$|\operatorname{dir}(T)\cdot V(p)|\leq\delta\quad\textrm{for all}\ (T,p)\in\mathbb{T}\times E_{\mathbb{T}}\ \textrm{with}\ p\in Y(T),$
@@ 6.2 Extremal collection of tubes, and multi-scale structure
$\Big{|}\bigcup_{T\in\mathbb{T}}Y(T)\Big{|}\leq\delta^{\sigma^{\prime}}.$
@@
Let $\varepsilon,\delta>0$. We say a pair $(\mathbb{T},Y)_{\delta}$ is $\varepsilon$-extremal if the following holds:
@@
    $\Big{|}\bigcup_{\mathbb{T}}Y(T)\Big{|}\leq\delta^{\sigma-\varepsilon}.$
It is immediate from the above definitions that for every $\varepsilon,\delta_{0}>0$, there exists an $\varepsilon$-extremal collection of tubes $(\mathbb{T},Y)_{\delta}$ for some $\delta\in(0,\delta_{0}]$; cf. [16, Lemma 3.1].
@@
For all $\varepsilon>0$, there exists $\eta>0$ and $\delta_{0}>0$ so that the following holds for all $\delta\in(0,\delta_{0}]$. Let $(\mathbb{T},Y)_{\delta}$ be an $\eta$-extremal collection of tubes, and let $\rho\in[\delta^{1-\varepsilon},\delta^{\varepsilon}]$. Then there is a refinement $(\mathbb{T}^{\prime},Y^{\prime})_{\delta}$ of $(\mathbb{T},Y)_{\delta}$ and a balanced cover $(\tilde{\mathbb{T}},\tilde{Y})_{\rho}$ of $(\mathbb{T}^{\prime},Y^{\prime})_{\delta}$ with the following properties:
@@
The proof is nearly identical to the proof of the corresponding statement in [16, Proposition 3.2]. In [16], the first major step was to establish the existence of a cover $\mathbb{T}_{\rho}$ of $\mathbb{T}$ (or more accurately, a large subset $\mathbb{T}^{\prime}\subset\mathbb{T}$), with the property that both $\mathbb{T}_{\rho}$ and the rescaled sets $\mathbb{T}^{\prime}[T_{\rho}]$ are sticky. In our case, this step was handled by Lemma 2.13. The second major step was to establish a shading on $\mathbb{T}_{\rho}$ and a refinement of the shadings on the sets $\mathbb{T}^{\prime}[T_{\rho}]$ so that not too many tubes pass through each point (and as a consequence, the corresponding collections are $\varepsilon$-extremal).
@@ 6.3 Local and global grains
For all $\varepsilon,\delta_{0}>0$, there exists $\delta\in(0,\delta_{0}]$ and an $\varepsilon$-extremal set of tubes $(\mathbb{T},Y)_{\delta}$ with the following two properties:
@@
    There is a 1-Lipschitz function $f:[-1,1]\rightarrow\mathbb{R}$ so that $(E_{\mathbb{T}}\cap\{z=z_{0}\})\cdot(1,f(z_{0}),0)$ is a $(\delta,1-\sigma,\delta^{-\varepsilon})_{1}$-ADset for each $z_{0}\in[-1,1]$.
@@
For all $\varepsilon,\delta_{0}>0$, there exists $\delta\in(0,\delta_{0}]$ and an $\varepsilon$-extremal set of tubes $(\mathbb{T},Y)_{\delta}$ with the following properties:
  1. 1
@@ 6.4 The slope function has large slope
For all $\varepsilon,\delta_{0}>0$, there exists $\delta\in(0,\delta_{0}]$ and an $\varepsilon$-extremal set of tubes $(\mathbb{T},Y)_{\delta}$ with the following properties:
  1. 1′′
@@
and that $(\mathbb{R}^{2}\times J)\cap\bigcup_{\mathbb{T}}Y(T)$ can be covered by a union of $\leq\delta^{-\eta}|J|^{-2+\sigma}$ cubes of side-length $J$.
@@
Here is a summary of Steps 1-3. After a refinement, we may suppose that for each point $p\in F:=(\mathbb{R}^{2}\times J)\cap\bigcup_{\mathbb{T}}Y(T)$,
@@
After a translation followed by a rotation that fixes the $z$ axis (i.e. a rotation of the form $(x,y,z)\mapsto(x\cos\theta-y\sin\theta,x\sin\theta+y\cos\theta,z)$), we may suppose that $J=[0,\rho^{1/2}]$ and $f(0)=0$.
Suppose for contradiction that $|f^{\prime}(z)|\leq|J|$ for some $z\in J$. Since $|f^{\prime\prime}(z)|\lesssim 1$, we have $|f^{\prime}(z)|\lesssim|J|$ for all $z\in J$, and thus $|f(z)|\lesssim|J|^{2}$. We find a subset $F_{2}\subset F$ with $|F_{2}|\geq\delta^{4\eta}|F|$ so that the follows holds:
@@
    has a large intersection with $F_{2}$, i.e.
@@
This finishes the summary of Steps 1-3 and we have introduced the necessary geometric objects and notation.
@@
But for each $T\in\mathbb{T}_{S}$, we have
@@
Recall that for each $T\in\mathbb{T}_{S}$, the (shaded) tube segment $F_{2}(T)$ is almost full. If $\tilde{V}(p_{S})$ was parallel to $\operatorname{dir}(T)$, then $F_{2}(T)\cdot\tilde{V}(p_{S})$ would fill out most of an interval of length $\rho^{1/2}$. However, (6.9) says that $F_{2}(T)\cdot\tilde{V}(p_{S})$ is an ADset with dimension $1-\sigma$. Since $\sigma>0$, we conclude that $\tilde{V}(p_{S})$ cannot be parallel to $\operatorname{dir}(T)$; in fact, the two vectors must be almost orthogonal. [16, Lemma 6.2] makes this precise; it says that if (6.9) holds, then
@@
$\Sigma_{S}=\{(x,y,z)\in[0,1]^{3}:|(x-x_{S},0,z)\cdot\tilde{V}(p_{S})|\leq\delta^{-2C_{0}\eta}\rho^{1/2}\}.$
@@
#$\#\mathbb{T}_{S}\leq|\Sigma_{S}|\delta^{-\eta}(\#\mathbb{T})\leq\delta^{-(2C+1)\eta}\rho^{1/2}(\#\mathbb{T}),$
@@
The proof of Proposition 6.5 is the same as the proof of [16, Proposition 6.1], except that we use Lemma 6.6 in place of [16, Lemma 6.3].
@@ 6.5 Twisted projections, and the completion of the proof
Recalling Definition 6.1 from [16], for $f\colon[-1,1]\to\mathbb{R}$ we define the twisted projection $\pi_{f}(x,y,z)=(x+f(z)y,z)$. The consequence of Proposition 6.5 is the following analogue of [16, Corollary 6.5]:
@@
For all $\varepsilon,\delta_{0}>0$, there exists $\delta\in(0,\delta_{0}]$; a pair $(\mathbb{T},Y)$ of $\delta$-tubes satisfying the Convex Wolff Axioms with error $\delta^{-\varepsilon}$ with a $\delta^{\varepsilon}$-dense shading; and a function $f\colon[-1,1]\to\mathbb{R}$ with $1\leq|f^{\prime}(z)|\leq 2$ and $|f^{\prime\prime}(z)|\leq 1/100$ for all $z\in[-1,1]$, so that
@@
To complete the proof of Theorem 5.2, it suffices to establish the following estimate, which is the analogue of [16, Proposition 7.1]:
@@
Comparing Lemma 6.7 and Proposition 6.8, we conclude that $\sigma=0$, which completes the proof of Theorem 5.2. It remains to prove Proposition 6.8. We defer this to the next section.
@@ 7 Twisted projections of tubes satisfying the Convex Wolff Axioms
Our goal in this section is to prove Proposition 6.8. In fact, we will prove a slightly stronger statement. We say a set of tubes $\mathbb{T}$ satisfies the Tube Wolff Axioms with error $C$, if for every $\rho$-tube $T_{\rho}$, at most $C|T_{\rho}|(\#\mathbb{T})$ tubes from $\mathbb{T}$ are contained in $T_{\rho}$; since $\rho$-tubes are convex, this is a special case of the Convex Wolff Axioms. We will prove the following mild generalization of Proposition 6.8:
@@
If $\mathbb{T}$ is a set of $\delta$-tubes contained in a bounded set such as $B(0,1)$, then we can identify the coaxial line of each tube with a point in $\mathbb{R}^{4}$. If $\mathbb{T}$ satisfies the Tube Wolff Axioms, then the corresponding set of points in $\mathbb{R}^{4}$ forms a $(\delta,2,O(1))$-Frostman set.
@@
where each $B_{i}$ in the covering is a cube and $r(B_{i})$ is the side-length of $B_{i}$. Next we recall Lemma 3.13 from [3]
Lemma 7.7.
Let $\delta,s>0$ and suppose that $\mathcal{H}^{s}_{\infty}(A)=\kappa>0$. Then there exists a $(\delta,s,100)$-Katz-Tao set $A^{\prime}\subset A$ with $\mathcal{E}_{\delta}(A^{\prime})\gtrsim\kappa\delta^{-s}$.
(In [3], the authors do not explicitly state the constant $C=100$, but this follows from their proof; any fixed constant would work equally well for our argument).
In light of Lemma 7.7, in order to prove Lemma 7.6, it suffices to obtain the bound
@@
Since $A$ is a $(\delta,s,C)$-Frostman set, $\mathcal{E}_{\delta}(A\cap B_{i})\leq C\lambda^{s}\mathcal{E}_{\delta}(A)$. The number of balls $B_{i}$ in the cover with $r(B_{i})\sim\lambda$ is $\gtrsim_{n,s}\lambda^{-s}|\log\delta|^{-1}C^{-1}$. This establishes (7.1) and concludes the proof. ∎
@@
Definition 7.8 ($\eta$-uniform set down to scale $\delta$).
@@
Let $\eta>0$ be a small parameter and $\delta\in 2^{-\eta^{-1}\mathbb{N}}$. Given an arbitrary set $A\subset[0,1]^{n}$, [13, Lemma 2.15] says there exists an $\eta$ uniform set $A^{\prime}\subset A$ down to scale $\delta$ and $\mathcal{E}_{\delta}(A^{\prime})\gtrsim|\log\delta|^{-\eta^{-1}}\mathcal{E}_{\delta}(A)$. We call $A^{\prime}$ an $\eta$-uniform refinement (usually $\delta$ is clear from context and we omit it from notation).
@@
Given an $\eta$-uniform set $A$ with $\delta$-covering number $\delta^{-s}$, the next lemma finds a Frostman-set if one zooms in at sufficiently small scale.
@@
Start with the output of [15, Lemma 5.21] applied to $A$. If there exists $t_{j}\geq t_{j+1}$, apply [15, Lemma 5.20] to merge the two intervals $[a_{j},a_{j+1}]$ and $[a_{j+1},a_{j+2}]$ to $[a_{j},a_{j+2}]$ and the two numbers $t_{j},t_{j+1}$ becomes $t_{j}^{\prime}$. Rename the set of intervals and $t_{j}$s and replace $J$ by $J-1$. Iterate the process until $t_{j}<t_{j+1}$ for all $j$. Now we have obtained a decomposition of scales $\{[a_{j},a_{j+1}]\}_{j=1}^{J}$ such that
@@
Divide the unit cube $[0,1]^{n}$ into a union of cubes $\tilde{Q}$ of side length $\rho^{s/d}\leq\rho$. For each $\tilde{Q}$, let $A_{\tilde{Q}}$ denote the translation that maps the center of $Q_{0}$ to the center of $\tilde{Q}$. Define $A=\bigsqcup A_{\tilde{Q}}(A_{0})$. Since the side-length of $\tilde{Q}$ is greater than the side-length of $Q_{0}$, $A_{\tilde{Q}}(A_{0})$ are disjoint, and so $\mathcal{E}_{\delta}(A)\gtrsim\delta^{-s}$.
@@
$\mathcal{E}_{\delta}(A\cap B_{r})\lesssim C\big{(}\frac{r}{\rho^{s/n}}\big{)}^{n}\cdot\big{(}\frac{\rho}{\delta}\big{)}^{s}\lesssim C\big{(}\frac{r}{\delta}\big{)}^{s}.\qed$
Next we will prove a weaker variant of Proposition 7.1, which relates the area of the twisted projection $\bigcup_{\mathbb{T}}\pi_{f}(Y^{\prime}(T))$ to the area of this projection at a coarser resolution.
@@
The quantity on the RHS of (7.2) will be closely related to the graph of cinematic functions (see [16, Definition 7.1]). To estimate it, we recall [16, Theorem 7.2]: